A ball rolls over the edge of a platform with only a horizontal velocity. The height of the platform is 1.0 m and the horizontal range of the ball from the base of the platform is 20 m. What is the horizontal velocity of the ball just before it touches the ground? Neglect air resistance.
Answer #1

Time taken by the ball to reach the ground

h=vyt+gt2/2

Since vy =0 ( initial vertical velocity ) ,h=1.0 m g=9.8 m/s2

1.0=0+9.8t2/2

1.0=4.9t2

t2=1/4.9=0.204082

t=0.452 s

Since no acceleration in the horizontal direction So

vx=x/t

vx=20/0.452

vx=44.25 m/s Answer